OpenStudy (anonymous):

If you exert a force of 700 N to walk 6 m up a flight of stairs in 6 s, how much power do you use?

OpenStudy (anonymous):

Power =work/time Work=f*d 700N*6m=4,200 4,200/6s=700j

OpenStudy (anonymous):

answer should be in Joules per second, since those are the units for power, energy (or work) over time.

OpenStudy (shamim):

Unit of power is also watt

OpenStudy (anonymous):

yes, Watt=Joules/seconds

OpenStudy (anonymous):

power = force * velocity velocity=d/t ie 6/6=1 power=700*1 = 700 Watt

OpenStudy (anonymous):

as power of moving objects is force into velocity

OpenStudy (anonymous):

just a note: @Ann-Hybrid even when you get the same value with your approach, it may not be totally correct, since you are assuming the movement happens at constant velocity and that is not stated in the problem description.